# Spiking Neural Network Architecture Search: A Survey

### v1

@@ III-F Non-differentiable Hardware Constraints
All of the search strategies mentioned above can be adapted to hardware-aware NAS (HW-NAS), which aims to optimize neural architectures to meet hardware-specific constraints such as memory, latency, and energy requirements. However, real-world hardware metrics such as bus widths, memory sizes, or specialized neuromorphic parameters are often discrete, making them non-differentiable. Although HW-NAS typically relies on differentiable loss functions for efficient optimization [93], these discrete choices complicate gradient-based approaches.
To address this challenge, methods like Gumbel-Softmax (used in FBNet [146]) introduce controlled noise to effectively “soften” discrete variables, while ProxylessNAS [145] either approximates the discrete search space with a continuous function or applies the REINFORCE algorithm [155] to handle binarized weights. These strategies facilitate backpropagation by approximating non-differentiable hardware constraints, allowing over-parameterized yet gradient-friendly models to incorporate crucial hardware considerations.
@@ III-G Performance Evaluation Metrics
Performance evaluation metrics are essential in SNNaS, as they guide the design of architectures that balance accuracy, latency, energy consumption, memory footprint, and area. While floating point operations (FLOPs) are commonly used, they can be misleading since equal FLOPs may yield varying latencies on different hardware [156, 135, 133]. Accuracy-focused but energy-aware NAS methods, such as MONAS [157] and NetAdapt [158], incorporate power consumption, and circuit area minimization has also been studied to reduce static power use [92]. For edge applications, limited memory budgets are critical, and latency or profiled energy often provide more reliable indicators than FLOPs alone. Balancing these hardware metrics alongside task performance ensures robust solutions for diverse deployment scenarios.
@@
To estimate hardware cost, NAS methods use analytical formulas for quick approximations, albeit with less precision on complex designs; lookup tables (LUTs) leverage pre-measured data (e.g., layer-specific latency and energy in [159]); and simulation-based approaches, such as MnasNet [139], offer detail at the expense of time. Machine learning models predict hardware behavior from architecture-hardware mappings [157, 160, 92, 161], but rely on high-quality data. Hybrid approaches combine these methods. For instance, fast analytical or LUT-driven screening may precede more rigorous simulations or ML estimates for top-performing candidates.
Table IV: Spiking Neural Network Architecture Search Papers
Model Search Space Search Algorithm Acceleration Strategy Performance Evaluation
Zhou et al. (2020) [13] Global Evolutionary Cheap Surrogate Accuracy
Nijenhuis (2021) [14] Cell-based Progressive NAS Weight Inheritance, Reduced Dataset, Lower Fidelity Estimates Accuracy
AutoSNN (2022) [7] Hierarchical Evolutionary One-Shot Accuracy and Energy Efficiency
SpikeDHS (2022) [15] Cell-based and Hierarchical Gradient-Based First-Order Gradient Approximation Accuracy, Sparsity, Computation Cost, Energy Consumption
GAQ (2022) [16] Global Evolutionary Accuracy and Memory Cost
SNASNet (2022) [8] Cell-based Random Search NAS without Training Accuracy
Auto-Spikformer (2023) [17] Global Evolutionary One-Shot Accuracy and Energy Consumption
NeuEvo (2023) [18] Hierarchical Evolutionary Accuracy and Sparsity
MixedSNN (2023) [19] Global Evolutionary NAS without Training Accuracy and PBSE
ANAS (2023) [77] Global Evolutionary Parallel Searching with Multi-Thread CanMore Simulator Energy-Delay (Latency and Energy Consumption)
Liu & Yi (2024) [20] Global Random Search NAS without Training GMG (Accuracy), Energy Consumption, Model Parameters
SpikeNAS (2024) [12] Cell-based Progressive NAS NAS without Training Accuracy and Memory Budget
Putra & Shafique (2024) [21] Cell-based Random Search NAS without Training Accuracy and Memory Footprint
HASNAS (2024) [22] Cell-based Grid Search NAS without Training Accuracy, Memory, Area, Latency, and Energy Consumption
MSE-NAS (2024) [23] Hierarchical Evolutionary NAS without Training BIE (Neural Activity)
Yan et al. (2024) [24] Global Gradient-Based One-Shot Accuracy and Computational Cost (SynOps)
LitE-SNN (2024) [25] Hierarchical Gradient-Based Weight Sharing Accuracy, Memory, Computational Cost (SynOps)
SpikeExplorer (2024) [26] Hierarchical Bayesian Optimization User-Set Constraints on Parameters Accuracy, Area, Latency, Energy Consumption
DESPINE (2024) [27] Hierarchical Evolutionary Accuracy, Energy Consumption (Spike Rate)
AutoST (2024) [28] Global Evolutionary NAS without Training FLOPs
Che et al. (2024) [82] Cell-based and Hierarchical Gradient-Based, DGS, TPS Limited Search Space, First-Order Gradient Approximation, Merging Preprocessing Accuracy, Sparsity, Computation Cost, Energy Consumption
SSTNAS (2024) [81] Cell-based Evolutionary NAS without Training Accuracy
ANCoEF (2024) [84] Hierarchical Reinforcement Learning One-Shot Accuracy, Hardware Performance, Energy Consumption, Area
@@ III-H Search Acceleration Strategies
A graph showing learning-curve extrapolation. The x-axis is ‘Training Epochs.’ The y-axis is ‘Performance Metric (e.g., Accuracy or Loss).’ A solid blue curve starts at the origin and rises quickly before bending toward a plateau, showing observed performance during the early part of training. A vertical gray line labeled ‘cutoff’ marks the last observed epoch. From the cutoff to the right, a dashed green curve continues upward with a gentle slope, representing the extrapolated (predicted) performance for later epochs. A legend in the lower right identifies the solid blue line as ‘Observed performance’ and the dashed green line as ‘Extrapolated performance.’ The figure illustrates using partial training data to predict final performance, allowing early stopping of models that are unlikely to improve significantly.
Other methods reduce full-training overhead even more drastically. One-shot NAS (weight sharing), exemplified by ENAS [76], DARTS [144], and SNAS [113], trains a supernet whose subnets share weights. In the SNN domain, AutoSNN [7], ANCoEF [84], Auto-Spikformer [17], and Yan et al.[24] adopt this strategy to accelerate architecture search without separate training for each candidate. Moving further, zero-shot approaches predict performance without training at all, using metrics like activation overlap[167, 168, 169, 170, 171]. In SNNaS, Liu et al. [20] estimate Loihi-based architectures using gradient-based Gram matrices, while Kim et al.[8] introduce the Sparsity-Aware Hamming Distance for zero-shot SNN designs. Follow-up works like SpikeNAS[12], HASNAS [22], and Li et al. [81] leverage similar metrics to achieve orders-of-magnitude faster searches, balancing computational efficiency with competitive accuracy.
@@ III-I Temporal Considerations
Temporal dynamics are central to SNNs’ unique capabilities, yet many SNNaS approaches adapt ANN-centric methods that overlook time-related factors. Recent work has begun to address this gap. For instance, LitE-SNN [25] integrates both spatial and temporal compression, and MixedSNN [19] leverages Period-Based Spike Evaluation (PBSE) to capture critical temporal attributes. Che et al. [82] further introduce a spatial-temporal search methodology that uses Temporal Parameter Search (TPS) to fine-tune different neurons’ time constants, optimizing both architecture and temporal behaviors. Such strategies signal an important transition toward a more holistic SNN design paradigm, embracing temporal as well as spatial dimensions to enhance neuromorphic computing.
@@ IV Challenges and Future Directions
Table IV outlines recent SNNaS studies surveyed in this work, highlighting their search spaces, search algorithms, acceleration strategies, and performance criteria. While these approaches have driven notable progress in the field, significant challenges remain regarding flexibility, interpretability, scalability, and broader applicability, especially beyond core tasks like image classification.
SNNaS faces persistent challenges regarding search space flexibility, interpretability, and reproducibility across diverse tasks. Human-designed or cell-based search spaces risk omitting genuinely novel architectures, while many methods remain tailored to specific datasets (often image classification) and do not naturally generalize to broader domains [172, 173]. Moreover, understanding why certain automatically discovered architectures perform well remains elusive, underscoring a need for deeper theoretical insights into SNN behaviors [174]. Resource-intensive search processes further restrict accessibility and reproducibility, despite partial solutions like one-shot NAS, weight sharing, or zero-cost proxies [73, 173]. These acceleration techniques can introduce biases—for instance, parameter sharing may degrade final model quality or misrank promising designs [175, 117, 176, 132].
@@
Simultaneously optimizing architectures and hyperparameters could lessen human bias and yield more robust solutions, especially under few-shot learning or meta-learning conditions [174, 177]. There is also growing interest in architectures resilient to adversarial attacks [73], multi-task learning, and multi-objective optimization that accounts for both performance and resource efficiency [73]. In parallel, research on hardware constraints in SNNaS has remained limited, often defaulting to von Neumann architectures despite memory walls and other performance bottlenecks [93]. Co-optimizing SNN architectures with emerging paradigms—such as in-memory computing, quantization, and pruning—could provide far more bio-plausible and efficient solutions [178, 93, 179]. Finally, most SNNaS methods still focus on a single, universal architecture rather than tailoring designs for specific hardware devices, suggesting future work in building device-targeted or co-explored solutions [93].
V Conclusion

### v2

@@ III-F Non-differentiable Hardware Constraints
All of the search strategies mentioned above can be adapted to HW-NAS, which aims to optimize neural architectures to meet hardware-specific constraints such as memory, latency, and energy requirements. However, real-world hardware metrics such as bus widths, memory sizes, or specialized neuromorphic parameters are often discrete, making them non-differentiable. Although HW-NAS typically relies on differentiable loss functions for efficient optimization [15], these discrete choices complicate gradient-based approaches.
To address this, methods like Gumbel-Softmax (used in FBNet [163]) introduce controlled noise to effectively “soften” discrete variables, while ProxylessNAS [28] either approximates the discrete search space with a continuous function or applies the REINFORCE algorithm [162] to handle binarized weights. These strategies facilitate backpropagation by approximating non-differentiable hardware constraints, allowing over-parameterized yet gradient-friendly models to incorporate crucial hardware considerations.
@@ III-G Performance Evaluation Metrics
Performance evaluation metrics are essential in SNNaS, as they guide the design of architectures that balance accuracy, latency, energy consumption, memory footprint, and area. While floating point operations (FLOPs) are commonly used, they can be misleading since equal FLOPs may yield varying latencies on different hardware [24, 181, 155]. Accuracy-focused but energy-aware NAS methods, such as Multi-Objective Neural Architectural Search (MONAS) [59] and NetAdapt [172], incorporate power consumption, and circuit area minimization has also been studied to reduce static power use [171]. For edge applications, limited memory budgets are critical, and latency or profiled energy often provides more reliable indicators than FLOPs alone.
Some SNNaS methods track spike-specific metrics such as the number of spikes (NoS), synaptic operations (SynOps), and Bit-SynOps as proxies for energy consumption. Spike count (NoS) is used by AutoSNN [97] and NeuEvo [132] as a proxy for power, as low spike counts reduce communication overhead; however, this metric does not account for the computational cost of the synapses themselves. ESNN [170] and LitE-SNN [82] argue that synaptic operations (SynOps), representing AC operations, are more comprehensive indicators of memory and computational resource demands. Structured Pruning for Spiking Neural Networks via Synaptic Operation Estimation and Reinforcement Learning (SPEAR) [165] found that after fine-tuning, SynOps change irregularly, suggesting that using pre-training SynOps as a hard constraint in NAS may result in final models that violate those constraints. Additionally, LitE-SNN [82] and GAQ-SNN [99] introduce mixed-precision quantization to NAS, arguing that hardware performance is determined by the “bit-width per operation" rather than just the number of operations. Balancing these hardware metrics with task performance ensures robust solutions for diverse scenarios.
@@
To estimate hardware cost, NAS methods use analytical formulas for quick approximations, albeit with less precision on complex designs; LUTs leverage pre-measured data (e.g., layer-specific latency and energy in [156]); and simulation-based approaches, such as mobile NAS (MNAS) [146], offer detail at the expense of time. Machine learning models predict hardware behavior from architecture-hardware mappings [59, 139, 171, 28], but rely on high-quality data. Hybrid approaches combine these methods. For instance, fast analytical or LUT-driven screening may precede more rigorous simulations or ML estimates for top-performing candidates.
While NAS benchmarks are well established for ANNs, SpikeNAS-Bench [142] represents the first benchmark explicitly designed for SNNaS. It provides a standardized, cell-based search space with spiking neuron operations and explicit temporal execution. Beyond accuracy, it adopts multi-objective evaluation metrics, including spike activity and synaptic operations, and supports proxy-based evaluation to reduce the cost of exhaustive SNN training. This benchmark enables controlled comparisons of NAS algorithms under consistent SNN-specific assumptions, addressing a key gap in prior SNNaS evaluation.
@@ III-H Search Acceleration Strategies
Search acceleration strategies in SNNaS aim to mitigate the prohibitive computational costs of training every candidate architecture from scratch [188, 119, 120] through approximations, learned parameter reuse, surrogate modeling, and hardware-aware pruning.
III-H1 Training-Fidelity Reduction Strategies
A common acceleration strategy evaluates architectures under reduced training fidelity, e.g., fewer epochs, truncated time steps, smaller datasets, or surrogate objectives. For example, learning curve extrapolation (LCE) [13] (illustrated in Figure 6) reduces computational costs by extrapolating final performance from partially trained models, avoiding full convergence during architecture evaluation. Progressive strategies accelerate search by incrementally growing architectures and only further training promising partial networks, reducing exploration cost but risking premature pruning.
@@ III-H1 Training-Fidelity Reduction Strategies
The covariance matrix adaptation evolution strategy (CMA-ES) [186], a surrogate-assisted model, mitigates this by using Gaussian processes to provide reliable fitness estimates even with reduced real-network evaluations, lowering the search cost while preserving ranking fidelity.
III-H2 Weight Reuse and Supernet-Based Acceleration
Weight reuse strategies accelerate convergence by avoiding training from scratch. HotNAS [64], a hot start method, initializes candidate models from a strong base model to speed early training. Network morphism and weight inheritance further reduce overhead by making incremental architectural changes while retaining trained weights [100].
More aggressive reuse occurs in one-shot or weight-sharing NAS, where a supernet is trained once, and candidate subnets inherit shared parameters. Prominent examples include Efficient NAS (ENAS) [109], DARTS [81], and SNAS [166], with SNN adaptations such as AutoSNN [97], ANCoEF [180], Auto-Spikformer [30], ESNN [170], and MA-DARTS [90]. ESNN [170], a one-shot NAS method, reduces search time by up to $67\times$ by using a branchless spiking supernet that eliminates the need for multiple candidate branches, lowering GPU memory consumption. While effective at reducing training costs, these methods are prone to ranking disorder and performance collapse, as shared weights often fail to reflect the true performance of independently trained SNNs, particularly under temporal coding schemes like time-to-first-spike (TTFS) [105, 170, 3].
III-H3 Initialization-Time and Zero-Shot Performance Proxies
Zero-shot approaches in the form of zero-cost (ZC) proxies rank architectures without gradient-based training by exploiting initialization-time statistics or analytically derived signals, such as activation overlap [94, 33, 169, 143, 87] or gradient-based measures [84]. They are well established in ANN-based NAS, but standard ANN proxies fail in the SNN domain because they rely on gradients that are inaccurate due to the non-differentiability of the spiking function [160].
SNN-specific proxies are explicitly designed to overcome these issues. For instance, SNASNet introduced Sparsity-Aware Hamming Distance (SAHD) [68] to normalize the batch-to-batch sparsity fluctuations in spiking neurons that skew distance-based architecture evaluations, enabling initialization-time, training-free (no weight optimization) scoring that fairly compares architectures under sparse spiking dynamics. Later, MONAS-ESNN proposed Adjusted SAHD (ASAHD) [125] to account for neuron-level sparsity variability in Parametric LIF (PLIF) neurons. While Liu et al. [84] estimate Loihi-based architectures using gradient-based Gram matrices (GMG), SpikeNAS [114] and NeuroNAS [111] use similar SNN-aware proxies to achieve orders-of-magnitude faster searches. For spiking transformers, AutoST [160] employs FLOPs as a gradient-independent proxy. While FLOPs are a poor global indicator of SNN efficiency, AutoST shows that within narrow architectural bands—where neuron models, temporal encoding schemes, and execution paradigms are fixed—FLOPs can act as a coarse discriminator between closely related candidate architectures. In these constrained regimes, FLOPs outperform spike-based proxies that are unstable under sparse, non-differentiable dynamics.
The MSE-NAS framework [104] introduces brain-inspired indirect evaluation (BIE) as a specialized "indirect evaluation" function. It accelerates the search by $60\times$ compared to direct training while ensuring transferability and robustness across datasets.
III-H4 Learned Performance Predictors and Few-Shot Methods
Bridging the gap between zero-shot speed and full-training reliability, several methods employ learned performance predictors trained on small subsets of evaluated architectures. EB-NAS [105] has an online predictor, CART (Classification and Regression Trees), that is updated dynamically during the search process to improve its accuracy as new architecture samples are trained, while PESNN [159] uses a Random Forest-based few-shot predictor to estimate accuracy at a reduced evaluation cost.
However, SNN predictors trained on limited samples struggle due to the highly non-linear temporal dynamics of SNNs, where minor architectural changes can induce abrupt changes in spike patterns. As a result, these predictors require costly seeding data and generalize poorly beyond the sampled design space, limiting their effectiveness in large-scale or hardware-constrained SNN searches [3]. The SPEAR framework [165] uses a few-shot predictor that employs linear regression to learn the relationship between pre-finetuning and post-finetuning synaptic operations, which solves a common problem of irregular SynOps changes during training.
Auxiliary learning mechanisms can also serve as acceleration tools. MA-DARTS [90] incorporates an auxiliary classifier to improve gradient propagation and stabilize training during differentiable search.
III-H5 Hardware-Aware and System-Level Acceleration
Frameworks like SPEAR [165] and NeuroNAS [111] accelerate searches by incorporating hardware constraints (e.g., memory, area, latency) directly into the search algorithm to prune the search space of unfeasible architectures before they are ever evaluated for accuracy.
The search for hardware-deployable SNNs is often limited by the speed of hardware evaluation. ANCoEF [180] accelerates the search loop by replacing standard simulators (like CanMore) with TrueAsync. TrueAsync uses Akka actor models to provide a $2\times$ to $15.8\times$ runtime speedup, directly addressing the high communication latency at the hardware/software boundary.
III-H6 Acceleration in Specialized SNN Paradigms (LSMs, TTFS)
In Liquid State Machine (LSM) NAS, CMA-ES [186] combined with a Gaussian Process (GP)-assisted surrogate significantly reduces the number of required evaluations. To mitigate supernet instability under TTFS constraints, TNAS-ER [83] pairs the spiking model with an identity-mapped ANN counterpart. While effective for stabilizing search, this strategy necessarily constrains the architecture to ANN-compatible representations and limits the exploitation of fully spiking temporal dynamics.
III-H7 Reliability Limits and Failure Modes of Acceleration in SNNaS
Despite their effectiveness, acceleration strategies in SNNaS face fundamental reliability constraints. Many zero-cost proxies depend on gradients or activation statistics that are ill-defined under sparse, non-differentiable spiking dynamics, leading to noisy or misleading rankings. Initialization sensitivity is especially severe for adaptive neurons such as PLIF, where sparsity evolves over time.
Weight-sharing methods often exacerbate ranking disorder, as shared weights fail to reflect the true performance of independently trained SNNs, especially under temporal coding schemes such as TTFS [83]. Similarly, predictor-based methods struggle to extrapolate beyond limited training samples due to the discontinuous nature of spike dynamics. As a result, acceleration methods transferred from ANN-NAS may bias the search toward architectures that score well on static or early-time proxies but underperform when full temporal dynamics and training convergence are considered [83].
@@ III-I Temporal Considerations
Temporal dynamics are central to SNNs’ unique capabilities, yet many SNNaS approaches adapt ANN-centric methods that overlook time-related factors. Recent work has begun to address this gap. For instance, LitE-SNN [82] integrates both spatial and temporal compression, and MixedSNN [167] leverages Period-Based Spike Evaluation (PBSE) to capture critical temporal attributes across the entire simulation duration. Che et al. [31] further introduce a spatial-temporal search methodology that uses TPS to fine-tune different neurons’ time constants, optimizing both architecture and temporal behaviors. MA-DARTS [90] incorporates temporal attention alongside spatial and channel dimensions, allowing the network to adaptively weight different time steps to focus on temporally important information while maintaining computational efficiency. By explicitly modeling temporal dynamics via attention mechanisms, MA-DARTS achieves significant improvements, suggesting that future SNNaS methods should integrate temporal processing into the architecture search process to support a more holistic SNN design paradigm.
Table IV: Global View of SNNaS Methods Grouped by Search Strategy. Search spaces, acceleration techniques, and metrics reflect dominant patterns within each category.
Strategy Representative Works Search Spaces Acceleration Key Metrics
Evolutionary EONS [128], MOO-SNNs [43], GP-Assisted CMA-ES [186], Spiking WANN [11], AutoSNN [97], GAQ [99], Auto-Spikformer [30], NeuEvo [132], MixedSNN [167], ANAS [179], ESL-SNNs [133], SPENSER [25], EB-NAS [105], MSE-NAS [104], DESPINE [5], AutoST [160], SSTNAS [74], PESNN [159], TNAS-ER [83], EMO-SNAS [138], MONAS-ESNN [125], Membrane-NAS [78], ABC-Arc [182], BCNAS-SNN [40], Vasilache et al. [153], NAS-TinyML [34] Global, Hierarchical, Sequential, Cell-based Training-free, One-shot, Surrogate models, Weight inheritance, Lower fidelity Accuracy, NoS, Energy, Network size
Gradient-Based Nijenhuis [100], SpikeDHS [29], SpikeDHS+TPS/HNS [31], ESNN [170], LitE-SNN [82], MA-DARTS [90], ST-DANO [178], SpikeX [168] Cell-based, Hierarchical, Global One-shot, Weight sharing, Lower fidelity Accuracy, SynOps, Task loss, EDP
Bayesian Opt. deepHyper [173], Benmeziane et al. [16], SpikeExplorer [102], ASNPC [157] Global, Cell-based Surrogate models, Lower fidelity, Weight inheritance Accuracy, Area, Latency, Power
RL-Based ANCoEF [180], SPEAR [165] Sequential One-shot, Lower fidelity Accuracy, EDP, SynOps, Area
Random/Grid SNASNet [68], Liu & Yi [84], SpiKernel [112], NeuroNAS [111], SpikeNAS [114], STMixer [42], LightSNN [3] Cell-based, Hierarchical, Sequential Training-free NAS, Lower fidelity, Progressive search SAHD, GMG, Accuracy, FLOPs
@@ IV Challenges and Future Directions
Table IV provides a global view of representative SNNaS studies surveyed in this work, highlighting their search spaces, search algorithms, acceleration strategies, and performance criteria. While these approaches have driven notable progress in the field, SNNaS continues to face fundamental challenges in search space design, hardware alignment, interpretability, and reproducibility.
IV-A Open Challenges in SNNaS
Many SNNaS methods remain misaligned with the realities of neuromorphic execution. Search objectives are typically evaluated on synchronous simulators and simplified energy proxies, overlooking asynchronous timing errors, communication effects, and non-numerical hardware design dimensions such as routing or on-chip dataflow. As a result, architectures that appear optimal in simulation may perform poorly when deployed. Moreover, understanding why certain automatically discovered architectures perform well remains elusive, underscoring a need for deeper theoretical insights into SNN behaviors [55].
Reproducibility represents a structural barrier to cumulative progress in SNNaS. Across NAS more broadly, unmaintained codebases, dependency decay, and inadequate documentation have led to a reproducibility crisis, with approximately 75% of published NAS methods failing to run out of the box [151]. SNNaS inherits these issues, which divert research effort toward reimplementation rather than extension, slowing cumulative advances.
These problems are exacerbated by the absence of large, diverse, and hardware-grounded SNNaS benchmarks, which continues to confine SNNaS to ANN-centric tasks, search spaces, and metrics. Although acceleration strategies such as one-shot NAS, weight sharing, and zero-cost proxies reduce computational burden [46, 161], they introduce additional biases: parameter sharing may degrade final model quality or misrank promising designs [70, 176, 183, 27].
Without stable implementations, standardized benchmarks, and environment-robust releases, the empirical validity of SNNaS comparisons remains fragile, particularly for hardware-aware and acceleration-heavy methods.
IV-B Emerging Research Directions
@@
Joint optimization of architectures and hyperparameters could lessen human bias and yield more robust solutions, especially under few-shot learning or meta-learning conditions [55, 77]. There is also growing interest in architectures resilient to adversarial attacks [46], multi-task learning, and multi-objective optimization that accounts for both performance and resource efficiency [46].
In parallel, hardware awareness in SNNaS remains limited. Many methods implicitly assume von Neumann execution models, overlooking memory bottlenecks and the constraints of neuromorphic systems [15]. Co-optimizing SNN architectures with emerging hardware-oriented paradigms such as in-memory computing, quantization, and pruning could enable more efficient and biologically plausible solutions [63, 15, 8]. Many existing SNNaS methods target a single “universal” architecture rather than exploring device-specific designs, highlighting the need for hardware-targeted or co-explored SNNaS frameworks [15].
Sustained progress in SNNaS also depends on treating reproducibility as a first-class research objective. Beyond open-source releases, future work should prioritize environment-stable implementations (e.g., containerized methods), clear documentation and pseudocode, explicit seed control and logging, and framework-level standardization to enable fair comparison and long-term reuse [151].
Finally, advancing SNNaS benchmarks to include diverse neuromorphic tasks, richer temporal dynamics, and hardware-grounded evaluation is critical. Such benchmarks should move beyond ANN-centric protocols to include SNN-native search spaces with explicit neuron-level parameterization, multi-objective metrics that capture accuracy, sparsity, and energy-relevant costs (e.g., SynOps), and standardized validation of acceleration proxies through rank correlation with fully trained models.
V Summary of Key Insights
This survey has examined the rapidly evolving field of Spiking Neural Network Architecture Search (SNNaS) through the lens of hardware-software co-design. Our analysis reveals that successful SNNaS fundamentally differs from traditional ANN-based NAS due to the unique computational characteristics of spiking networks. The non-differentiable spike function, temporal dynamics spanning multiple timesteps, and event-driven processing paradigm collectively necessitate specialized search spaces, evaluation metrics, and optimization strategies (Section II). A recurring theme across the surveyed literature is that approaches that simply adapt ANN architectures, e.g., replacing ReLU with LIF neurons or borrowing backbone structures like VGG and ResNet, consistently outperform methods designed specifically for spike-based computation (Section III-A). The evolution from global and sequential search spaces toward cell-based, hierarchical, and biologically-inspired representations reflects the field’s growing recognition that SNN-specific inductive biases are essential for discovering high-performing architectures (Section III-B).
The co-design perspective emerges as perhaps the most critical insight from this survey. Hardware constraints in neuromorphic computing are not merely optimization targets to be addressed post-hoc; they fundamentally shape which architectures are viable and efficient (Section III-C). Asynchronous, event-driven platforms like Intel’s Loihi achieve their power efficiency by eliminating global clocks, but this design choice renders certain operations problematic. For instance, spiking matrix multiplication and max-pooling layers become vulnerable to timing imprecision (Section III-D2). Co-exploration frameworks that jointly search neural architectures and hardware configurations (e.g., ANAS, ANCoEF) consistently achieve superior energy-delay products compared to methods that optimize these dimensions separately (Section III-D1). Furthermore, the choice of performance metrics profoundly impacts search outcomes. While spike count (NoS) serves as a simple proxy for power consumption, synaptic operations (SynOps) provide more comprehensive estimates of computational cost, and Bit-SynOps account for the impact of quantization on hardware efficiency (Section III-G).
Finally, the field has made remarkable progress in addressing the computational burden of architecture search through training-free and low-fidelity evaluation strategies (Section III-H). Zero-shot metrics like SAHD and GMG enable architecture ranking without any training, while one-shot supernet approaches and surrogate predictors dramatically reduce the cost of evaluating candidate architectures. These acceleration techniques have reduced search times from thousands of GPU-hours to practical levels, democratizing SNNaS research. Additionally, the explicit modeling of temporal dynamics, through temporal attention mechanisms, timestep optimization, and spike-timing-aware evaluation, represents a critical frontier that remains underexploited in current methods (Section III-I). As summarized in Table IV, the diversity of search strategies (evolutionary, gradient-based, Bayesian, and reinforcement learning approaches) reflects the field’s ongoing exploration of how best to navigate the complex, multi-objective optimization landscape inherent to SNNaS (Section III-E). Looking forward, significant challenges remain in interpretability, generalization beyond image classification, and scalability (Section IV). As neuromorphic hardware continues to mature and standardized tools emerge, we anticipate that SNNaS will play an increasingly central role in unlocking the full potential of brain-inspired computing for energy-efficient, real-time AI applications.
VI Conclusion